Verify that the given functions form the fundamental set of solution of the differential equation on the given indicated interval.


2x^2y" +5xy' + y = x^2-x

y = C1x^-1/2 + C2x^-1 + 15x^2-1/6x, (-0, infinity)


1
Expert's answer
2021-10-25T16:34:37-0400
"2x^2y'' +5xy' + y = x^2-x"

"y = C_1x^{-1/2} + C_2x^{-1} +\dfrac{1}{15}x^2-\dfrac{1}{6}x, (0, \infin)"

"y'=-\dfrac{1}{2}C_1x^{-3/2} - C_2x^{-2} +\dfrac{2}{15}x-\dfrac{1}{6}"

"y''=\dfrac{3}{4}C_1x^{-5/2} +2 C_2x^{-3} +\dfrac{2}{15}"

"2x^2y'' +5xy' + y = x^2-x"

"2x^2(\dfrac{3}{4}C_1x^{-5/2} +2 C_2x^{-3} +\dfrac{2}{15})"

"+5x(-\dfrac{1}{2}C_1x^{-3/2} - C_2x^{-2} +\dfrac{2}{15}x-\dfrac{1}{6})"

"+C_1x^{-1/2} + C_2x^{-1} +\dfrac{1}{15}x^2-\dfrac{1}{6}x"

"=\dfrac{3}{2}C_1x^{-1/2} +4 C_2x^{-1} +\dfrac{4}{15}x^2"

"-\dfrac{5}{2}C_1x^{-1/2} -5 C_2x^{-1} +\dfrac{2}{3}x^2-\dfrac{5}{6}x"

"+C_1x^{-1/2} + C_2x^{-1} +\dfrac{1}{15}x^2-\dfrac{1}{6}x"

"=x^2-x, True\ x\in(0, \infin)"